# On the contraction properties of Sinkhorn semigroups

### v2

@@ 5.6 Exponential marginal models
Note that the functions $x\in[0,\infty[\mapsto\tau_{\iota}x\in[0,\infty[$ with $\iota\in\{u,v\}$ have compact sub-level sets. Let ${\cal Q}(x,dy)$ be the exponential transition defined for any $x\in[0,\infty[$ by
@@
In this context, condition ${\cal H}^{\prime}_{\delta}(U,V)$ holds for any $({\varsigma}/{\tau_{u}})\vee({\varsigma}/{\tau_{v}})<\delta<{1}/{2}$.
5.7 Semi-compact models
@@
Semi-discrete model
@@
Let $\nu_{V}(dy)=e^{-V(y)}~dy$ be a Boltzmann-Gibbs probability measure on $\mathbb{R}^{d}$ and assume that $(U,V,W)$ are continuous functions satisfying the polynomial growth conditions (83). In this context, the sub-level sets of $U$ are finite, thus compact in $\mathbb{X}$.
Following the proof of Proposition 5.1 these conditions ensure that ${\cal H}^{\prime}_{\delta}(U,V)$ holds for some $0<\delta<1/2$ as soon as condition (84) is satisfied. The analysis of heavy tailed transitions follows word-for-word the same lines of arguments as in Section 5.2, this it is skipped.
Semi-compact model
@@
for some continuous function $w(x)$ and some parameters $a\leq 2$ and $p\leq 1$. These models satisfy ${\cal H}^{\prime}_{\delta}(U,V)$ as soon as
$(p=2~~\mbox{\rm and}~~b<\delta/2)\quad\mbox{\rm or}\quad(p<2)$ (86)
@@
5.8 Linear-Gaussian models
Consider the state spaces $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, for some integer $d\geq 1$ and the Lebesgue measure $\lambda(dz)=\nu(dz)=dz$. Hereafter, points $x$ in the Euclidean space $\mathbb{X}=\mathbb{R}^{d}$ are represented by $d$-dimensional column vectors (or, equivalently, by $d\times 1$ matrices). When there is no possible confusion, we use the notation $\|\cdot\|$ for any equivalent matrix or vector norm.
Let ${\cal G}l_{d}$ denote the general linear group of $(d\times d)$-invertible matrices, and ${\cal S}^{+}_{d}\subset{\cal G}l_{d}$ the subset of positive definite matrices. We sometimes use the Löwner partial ordering notation $v_{1}\geq v_{2}$ to mean that a symmetric matrix $v_{1}-v_{2}$ is positive semi-definite (equivalently, $v_{2}-v_{1}$ is negative semi-definite), and $v_{1}>v_{2}$ when $v_{1}-v_{2}$ is positive definite (equivalently, $v_{2}-v_{1}$ is negative definite). Given $v\in{\cal S}_{d}^{+}$ we denote by $v^{1/2}$ the principal (unique) symmetric square root.
Denote by $\nu_{m,\sigma}$ the Gaussian distribution on $\mathbb{R}^{d}$ with mean $m\in\mathbb{R}^{d}$ and covariance matrix $\sigma\in{\cal S}^{+}_{d}$. In addition, let $g_{\sigma}$ denote the probability density function of the distribution $\nu_{0,\sigma}$, with covariance matrix $\sigma\in{\cal S}^{+}_{d}$. Consider the target probability measures $(\lambda_{U},\nu_{V})=(\nu_{m,\sigma},\nu_{\overline{m},\overline{\sigma}})$ for some given parameters
@@ 5.8 Linear-Gaussian models
$U(x)=-\log{g_{\sigma}(x-m)}\quad\mbox{\rm and}\quad V(y):=-\log{g_{\overline{\sigma}}(y-\overline{m})}.$ (87)
The linear Gaussian transition $\cal Q$} is defined by (2), with the potential
$W(x,y)=-\log{g_{\tau}(y-(\alpha+\beta x))}\quad$ (88)
@@
$\|\pi_{2n}-\nu_{\overline{m},\overline{\sigma}}\|_{\sf tv}\leq c~\rho^{n}\quad\mbox{\rm with}\quad\rho=(1+\lambda_{\sf min}(\Delta_{\varpi}))^{-1}<1.$ (89)
The matrix $\Delta_{\varpi}$ in expression (89) above is defined as
$\Delta_{\varpi}:=\frac{\varpi}{2}+\left(\varpi+\left(\frac{\varpi}{2}\right)^{2}\right)^{1/2}\geq\varpi+(I+\varpi^{-1})^{-1}$ (90)
@@
Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is satisfied with the Lyapunov functions
$\varphi(x):=1+(x-m)^{\prime}\sigma^{-1}(x-m)\quad\mbox{\rm and}\quad\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$ (91)
@@
The choice of the Lyapunov functions is not unique. For instance, Condition ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is also met for any $\delta>0$ with the exponential Lyapunov functions
@@
$\psi(y)$ $:=$ $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$ (92)
There also exists some $\delta_{0}>0$ such that ${\cal D}_{\varphi,\psi}({\cal S}^{\circ})$ is met for any $0<\delta<\delta_{0}$ with the exponential Lyapunov functions
@@
$\psi(y)$ $:=$ $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$ (93)
@@
We end this section with an equivalent formulation of condition ${\cal H}_{\delta}(U,V)$ in terms of the parameters of the Gaussian potentials (87) and (88).
@@
Condition ${\cal H}_{\delta}(U,V)$ is satisfied for the Lyapunov functions in (93) if, and only if,
$\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~\sigma^{-1}\quad\text{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~\overline{\sigma}^{-1}.$ (94)
@@
5.9 Finite mixture models
@@
$U=-\log{\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}}}\quad\mbox{\rm and}\quad V=-\log{\sum_{j\in E_{1}}\xi_{1}(j)~e^{-V_{j}}}.$ (95)
@@
Condition ${\cal H}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1[$ if conditions ${\cal H}_{\delta}(U_{i},V_{j})$ are met for the same $\delta\in]0,1[$ and almost every $(i,j)\in(E_{0}\times E_{1})$.
@@
Let us now consider the mixture model (95) with $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, the potential $W(x,y)$ defined in Eq. (88) and the log-densities
$U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$ (96)
where $(m_{i},\overline{m}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in(\mathbb{R}^{d}\times\mathbb{R}^{d})^{E_{0}\times E_{1}}$ and $(\sigma_{i},\overline{\sigma}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in({\cal S}^{+}_{d}\times{\cal S}^{+}_{d})^{E_{0}\times E_{1}}$ are given parameters. Combining Proposition 5.4 with Proposition 5.3 we readily obtain the following result.
@@
The Gaussian mixture model defined in (95) with the Gaussian log-densities of (96) defined on $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$ satisfies Condition ${\cal H}_{\delta}(U,V)$ for some $\delta\in]0,1[$ provided that
$\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~\sigma_{i}^{-1}\quad\mbox{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~\overline{\sigma}_{j}^{-1}$ (97)
@@
$\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~dx$ (98)
is a well known Gaussian mixture approximation of the distribution $\eta(dx)$ (as the number of samples $N\rightarrow\infty$). The covariance matrix $\sigma$ is a bandwidth parameter that controls the amount and orientation of the smoothing. Note that (98) coincides with (95) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$. For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [parzen, rosenblatt] as well as the books [chacon, simonov, wand].
The Gaussian mixture distribution (98) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [doucet-bortoli] and references therein.
@@
Schrödinger bridges and the Sinkhorn algorithm are often used in practice to solve high dimensional optimal transport problems. In our context, choosing $\tau=tI$ and $(\alpha,\beta)=(0,I)$ in (88), the Schrödinger bridge in the limit $t\rightarrow 0$ of vanishing regularization solves the Monge-Kantorovich problem and converge towards the optimal $2$-Wasserstein coupling. In this case, ${\cal S}_{0}=Q$ coincides with the heat kernel on a time horizon $t$ (see also Remark 2.3 in [adm-24]), namely,
${\cal Q}(x,dy)=(2\pi t)^{-d/2}~\exp{\left(-\frac{1}{2t}~\|y-x\|^{2}\right)}~dy.$
In this context, the l.h.s. condition in (97) applied to (98) becomes
$\frac{1}{t}~I<(\delta\wedge(1-\delta))~\sigma^{-1}.$ (99)
In many cases, the Markov transition ${\cal Q}(x,dy)$ of interest is generated by a stochastic diffusion between time 0 and a time horizon $t$ –see, e.g., [leonard, stromme2023sampling, kato2024large]. In our context, if we choose $(\alpha,\beta)=(0,e^{tA})$ and $\tau=\tau_{t}:=\int_{0}^{t}e^{sA}e^{sA^{\prime}}ds$ for some Hurwitz matrix $A$ and some $t>0$, ${\cal Q}(x,dy)$ coincides with the transition semigroup of the Ornstein-Uhlenbeck diffusion (see Remark 2.3 in [adm-24]). In this case, the l.h.s. condition in (97) applied to (98) is met for any $t\geq t_{0}$ provided that
$e^{tA^{\prime}}\tau_{t_{0}}^{-1}e^{tA}<(\delta\wedge(1-\delta))~\sigma^{-1}.$ (100)
Both transitions gradually transform the data distribution $\lambda_{U}$ into a simpler distribution $\lambda_{U}{\cal Q}$ by injecting noise into data. Also note that (99) and (100) are met when the time horizon is chosen sufficiently large.
The choice of the prior distribution $\nu_{V}$ is rather flexible. For instance, we can choose a Gaussian distribution, say $\nu_{0,\overline{\sigma}}$, satisfying the r.h.s. condition in (94). Starting from such Gaussian random states, the backward Sinkhorn transitions (indexed by odd indices) eventually generates samples back into the initial distribution by slowly removing the noise distortion. Conditions (99) and (100) reflect the fact that the forward noise injection as well as the reverse denoising procedure in diffusion-based generative models often requires very large time horizons.
@@ Acknowledgements
J. Miguez acknowledges the financial support of the Office of Naval Research (award no. N00014-22-1-2647) and MICIU/AEI/10.13039/501100011033/FEDER UE (grant no. PID2021-125159NB-I00 TYCHE).

### v3

@@ 5.6. Exponential marginal models
Note that the functions $x\in[0,\infty[\mapsto\tau_{\iota}x\in[0,\infty[$ with $\iota\in\{u,v\}$ have compact sub-level sets. Let ${\mathcal{Q}}(x,dy)$ be the exponential transition defined for any $x\in[0,\infty[$ by
@@
In this context, condition ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ holds for any $({\varsigma}/{\tau_{u}})\vee({\varsigma}/{\tau_{v}})<\delta<{1}/{2}$.
5.7. Semi-compact models
@@
Semi-discrete model.
@@
Let $\nu_{V}(dy)=e^{-V(y)}~dy$ be a Boltzmann-Gibbs probability measure on $\mathbb{R}^{d}$ and assume that $(U,V,W)$ are continuous functions satisfying the polynomial growth conditions (5.2). In this context, the sub-level sets of $U$ are finite, thus compact in $\mathbb{X}$.
Following the proof of Proposition 5.1 these conditions ensure that ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ holds for some $0<\delta<1/2$ as soon as condition (5.3) is satisfied. The analysis of heavy tailed transitions follows word-for-word the same lines of arguments as in Section 5.2, this it is skipped.
Semi-compact model.
@@
for some continuous function $w(x)$ and some parameters $a\leq 2$ and $p\leq 1$. These models satisfy ${\mathcal{H}}^{\prime}_{\delta}(U,V)$ as soon as
(5.5) $(p=2~~\mbox{\rm and}~~b<\delta/2)\quad\mbox{\rm or}\quad(p<2)$
@@
5.8. Linear-Gaussian models
Consider the state spaces $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, for some integer $d\geq 1$ and the Lebesgue measure $\lambda(dz)=\nu(dz)=dz$. Hereafter, points $x$ in the Euclidean space $\mathbb{X}=\mathbb{R}^{d}$ are represented by $d$-dimensional column vectors (or, equivalently, by $d\times 1$ matrices).
Let ${\mathcal{G}}l_{d}$ denote the general linear group of $(d\times d)$-invertible matrices, and ${\mathcal{S}}^{+}_{d}\subset{\mathcal{G}}l_{d}$ the subset of positive definite matrices. We sometimes use the Löwner partial ordering notation $v_{1}\geq v_{2}$ to mean that a symmetric matrix $v_{1}-v_{2}$ is positive semi-definite (equivalently, $v_{2}-v_{1}$ is negative semi-definite), and $v_{1}>v_{2}$ when $v_{1}-v_{2}$ is positive definite (equivalently, $v_{2}-v_{1}$ is negative definite). Given $v\in{\mathcal{S}}_{d}^{+}$ we denote by $v^{1/2}$ the principal (unique) symmetric square root.
Denote by $\nu_{m,\sigma}$ the Gaussian distribution on $\mathbb{R}^{d}$ with mean $m\in\mathbb{R}^{d}$ and covariance matrix $\sigma\in{\mathcal{S}}^{+}_{d}$. In addition, let $g_{\sigma}$ denote the probability density function of the distribution $\nu_{0,\sigma}$, with covariance matrix $\sigma\in{\mathcal{S}}^{+}_{d}$. Consider the target probability measures $(\lambda_{U},\nu_{V})=(\nu_{m,\sigma},\nu_{\overline{m},\overline{\sigma}})$ for some given parameters
@@ 5.8. Linear-Gaussian models
(5.6) $U(x)=-\log{g_{\sigma}(x-m)}\quad\mbox{\rm and}\quad V(y):=-\log{g_{\overline{\sigma}}(y-\overline{m})}.$
The linear Gaussian transition $\mathcal{Q}$} is defined by (1.2), with the potential
(5.7) $W(x,y)=-\log{g_{\tau}(y-(\alpha+\beta x))}\quad$
@@
(5.8) $\|\pi_{2n}-\nu_{\overline{m},\overline{\sigma}}\|_{\sf tv}\leq c~\rho^{n}\quad\mbox{\rm with}\quad\rho=(1+\lambda_{\sf min}(\Delta_{\varpi}))^{-1}<1.$
The matrix $\Delta_{\varpi}$ in expression (5.8) above is defined as
(5.9) $\Delta_{\varpi}:=\frac{\varpi}{2}+\left(\varpi+\left(\frac{\varpi}{2}\right)^{2}\right)^{1/2}\geq\varpi+(I+\varpi^{-1})^{-1}$
@@
Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is satisfied with the Lyapunov functions
(5.10) $\varphi(x):=1+(x-m)^{\prime}\sigma^{-1}(x-m)\quad\mbox{\rm and}\quad\psi(y):=1+(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m}).$
@@
The choice of the Lyapunov functions is not unique. For instance, Condition ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is also met for any $\delta>0$ with the exponential Lyapunov functions
@@
(5.11) $\psi(y)$ $:=$ $\exp{\left(\delta\,((y-\overline{m})^{\prime}\,\overline{\sigma}^{-1}(y-\overline{m}))^{1/2}\right)}.$
There also exists some $\delta_{0}>0$ such that ${\mathcal{D}}_{\varphi,\psi}({\mathcal{S}}^{\circ})$ is met for any $0<\delta<\delta_{0}$ with the exponential Lyapunov functions
@@
(5.12) $\psi(y)$ $:=$ $\exp{\left(\delta\,(y-\overline{m})^{\prime}\overline{\sigma}^{-1}(y-\overline{m})\right)}.$
@@
We end this section with an equivalent formulation of condition ${\mathcal{H}}_{\delta}(U,V)$ in terms of the parameters of the Gaussian potentials (5.6) and (5.7).
@@
Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for the Lyapunov functions in (5.12) if, and only if,
(5.13) $\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~\sigma^{-1}\quad\text{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~\overline{\sigma}^{-1}.$
@@
5.9. Finite mixture models
@@
(5.14) $U=-\log{\sum_{i\in E_{0}}\xi_{0}(i)~e^{-U_{i}}}\quad\mbox{\rm and}\quad V=-\log{\sum_{j\in E_{1}}\xi_{1}(j)~e^{-V_{j}}}.$
@@
Condition ${\mathcal{H}}_{\delta}(U,V)$ is satisfied for some $\delta\in]0,1[$ if conditions ${\mathcal{H}}_{\delta}(U_{i},V_{j})$ are met for the same $\delta\in]0,1[$ and almost every $(i,j)\in(E_{0}\times E_{1})$.
@@
Let us now consider the mixture model (5.14) with $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$, the potential $W(x,y)$ defined in Eq. (5.7) and the log-densities
(5.15) $U_{i}(x)=\log{g_{\sigma_{i}}(x-m_{i})}\quad\mbox{\rm and}\quad-V_{j}(y):=\log{g_{\overline{\sigma}_{j}}(y-\overline{m}_{j})},$
where $(m_{i},\overline{m}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in(\mathbb{R}^{d}\times\mathbb{R}^{d})^{E_{0}\times E_{1}}$ and $(\sigma_{i},\overline{\sigma}_{j})_{(i,j)\in(E_{0}\times E_{0})}\in({\mathcal{S}}^{+}_{d}\times{\mathcal{S}}^{+}_{d})^{E_{0}\times E_{1}}$ are given parameters. Combining Proposition 5.4 with Proposition 5.3 we readily obtain the following result.
@@
The Gaussian mixture model defined in (5.14) with the Gaussian log-densities of (5.15) defined on $\mathbb{X}=\mathbb{Y}=\mathbb{R}^{d}$ satisfies Condition ${\mathcal{H}}_{\delta}(U,V)$ for some $\delta\in]0,1[$ provided that
(5.16) $\beta^{\prime}\tau^{-1}\beta<(\delta\wedge(1-\delta))~\sigma_{i}^{-1}\quad\mbox{and}\quad\tau^{-1}<(\delta\wedge(1-\delta))~\overline{\sigma}_{j}^{-1}$
@@
(5.17) $\lambda_{U}(dx):=\frac{1}{N}\sum_{1\leq i\leq N}g_{\sigma}(x-X_{i})~dx$
is a well known Gaussian mixture approximation of the distribution $\eta(dx)$ (as the number of samples $N\rightarrow\infty$). The covariance matrix $\sigma$ is a bandwidth parameter that controls the amount and orientation of the smoothing. Note that (5.17) coincides with (5.14) when $(m_{i},\sigma_{i})=(X_{i},\sigma)$ with the counting measure $\xi_{0}(i)=1/N$ on the index set $E_{0}=\{1,\ldots,N\}$. For a more detailed discussion on these non parametric statistical estimation techniques we refer to the pioneering articles [parzen, rosenblatt] as well as the books [chacon, simonov, wand].
The Gaussian mixture distribution (5.17) can be used to estimate complex and unknown data distributions arising in generative models, see for instance [doucet-bortoli] and references therein.
@@
Schrödinger bridges and the Sinkhorn algorithm are often used in practice to solve high dimensional optimal transport problems. In our context, choosing $\tau=tI$ and $(\alpha,\beta)=(0,I)$ in (5.7), the Schrödinger bridge in the limit $t\rightarrow 0$ of vanishing regularization solves the Monge-Kantorovich problem and converge towards the optimal $2$-Wasserstein coupling. In this case, ${\mathcal{S}}_{0}=Q$ coincides with the heat kernel on a time horizon $t$ (see also Remark 2.3 in [adm-24]), namely,
${\mathcal{Q}}(x,dy)=(2\pi t)^{-d/2}~\exp{\left(-\frac{1}{2t}~\|y-x\|^{2}\right)}~dy.$
with the Euclidian norm $\|z\|^{2}=z^{\prime}z$. In this context, the l.h.s. condition in (5.16) applied to (5.17) becomes
(5.18) $\frac{1}{t}~I<(\delta\wedge(1-\delta))~\sigma^{-1}.$
In many cases, the Markov transition ${\mathcal{Q}}(x,dy)$ of interest is generated by a stochastic diffusion between time 0 and a time horizon $t$ –see, e.g., [leonard, stromme2023sampling, kato2024large]. In our context, if we choose $(\alpha,\beta)=(0,e^{tA})$ and $\tau=\tau_{t}:=\int_{0}^{t}e^{sA}e^{sA^{\prime}}ds$ for some Hurwitz matrix $A$ and some $t>0$, ${\mathcal{Q}}(x,dy)$ coincides with the transition semigroup of the Ornstein-Uhlenbeck diffusion (see Remark 2.3 in [adm-24]). In this case, the l.h.s. condition in (5.16) applied to (5.17) is met for any $t\geq t_{0}$ provided that
(5.19) $e^{tA^{\prime}}\tau_{t_{0}}^{-1}e^{tA}<(\delta\wedge(1-\delta))~\sigma^{-1}.$
Both transitions gradually transform the data distribution $\lambda_{U}$ into a simpler distribution $\lambda_{U}{\mathcal{Q}}$ by injecting noise into the data. Also note that (5.18) and (5.19) are met when the time horizon is chosen sufficiently large.
The choice of the prior distribution $\nu_{V}$ is rather flexible. For instance, we can choose a Gaussian distribution, say $\nu_{0,\overline{\sigma}}$, satisfying the r.h.s. condition in (5.13). Starting from such Gaussian random states, the backward Sinkhorn transitions (indexed by odd indices) eventually generate samples back into the initial distribution by slowly removing the noise distortion. Conditions (5.18) and (5.19) reflect the fact that the forward noise injection as well as the reverse denoising procedure in diffusion-based generative models often require very large time horizons.
@@ Acknowledgements
We would like to thank the anonymous reviewers for their excellent suggestions for improving the paper. Their detailed comments greatly improved the presentation of the article.
@@
J. Miguez acknowledges the financial support of the Office of Naval Research (award no. N00014-22-1-2647), MICIU/AEI/10.13039/501100011033/FEDER UE (grants no. PID2021-125159NB-I00 TYCHE and PID2024-158181NB-I00 NISA), and Community of Madrid, under grant IDEA-CM (TEC- 2024/COM-89).
Appendix
This appendix compiles technical material and supplementary results that bolster the paper’s central arguments. Appendix A focuses on Gaussian Sinkhorn algorithms. This material is taken from [adm-24]. We review some properties of the Riccati matrix differential equations and the random maps associated with Sinkhorn transitions. We also provide equivalent formulations of the drift-minorization conditions for the Lyapunov functions presented in Proposition 5.2. Appendix B is devoted to basic properties of Hilbert’s projective metric. Appendix C recalls several analytical tools related to weighted Banach spaces, operator norms, and contraction coefficients that are used throughout the stability analysis. For a more thorough discussion on weighted norms and contraction coefficients, we refer the reader to the articles [dg-01, dm-03, horton-dp, delgerber2025] as well as the book [penev].
Appendix D reviews Sinkhorn potentials. The first part, Appendix D.1, presents some commutation formulae. Building on this framework, several quantitative estimates are provided in Appendix D.2. Section D.3 is dedicated to Lyapunov inequalities and minorization conditions. The Proof of Theorem 4.8 is provided in Appendix D.4 and Appendix E contains further technical arguments.